OpenStudy (anonymous):

I think the answer is 15 but I am not sure because I have no idea what the ! means. So if someone could help that would be great. What is the value of the expression? 5!(8 – 5)!

OpenStudy (unanimoose):

Basically look at it this way: \[5! = 5 \times 4 \times 3 \times 2 \times 1\]

OpenStudy (unanimoose):

Now in (8 - 5) you have to do what is inside the parenthesis before you actuallty determine to factorial.

OpenStudy (anonymous):

Oh so the ! means that number times all the numbers that are smaller?

OpenStudy (unanimoose):

you're not gonna do 8! and 5! because that would be a long answer.

OpenStudy (unanimoose):

Yes.

OpenStudy (anonymous):

Thank you

OpenStudy (unanimoose):

So (8-5)! is 3! not 8! - 5!

OpenStudy (unanimoose):

So first do: 5 * 4 * 3 * 2 3 * 2 You don't have to put in the 1 because that really doesn't do anything. Once you did that you just add em up and you got the answer. Let me know what you get. :)

OpenStudy (anonymous):

126

OpenStudy (unanimoose):

Wait no, multiply don't add. My bad. 5!(3)!

OpenStudy (unanimoose):

So 120 * 6

OpenStudy (anonymous):

720

OpenStudy (unanimoose):

Yep. :P

OpenStudy (anonymous):

Thank you @Unanimoose

OpenStudy (unanimoose):

Np. :)